# P3-LLM: An Integrated NPU-PIM Accelerator for LLM Inference Using Hybrid Numerical Formats

### v1

@@ Abstract
The substantial memory bandwidth and computational demand of large language models (LLMs) present critical challenges for efficient inference. To tackle this, the literature has explored heterogeneous systems that combine neural processing units (NPUs) with DRAM-based processing-in-memory (PIM) for LLM acceleration. However, existing high-precision (e.g., FP16) PIM compute units incur significant area and power overhead in DRAM technology, limiting the effective computation throughput. In this paper, we introduce P3-LLM, a novel NPU-PIM integrated accelerator for LLM inference using hybrid numerical formats. Our approach is threefold: First, we propose a flexible mixed-precision quantization scheme, which leverages hybrid numerical formats to quantize different LLM operands with high compression efficiency and minimal accuracy loss. Second, we architect an efficient PIM accelerator co-design for P3-LLM, featuring lightweight compute units to support our hybrid numerical formats. The enhanced PIM compute units significantly boost the computation throughput under iso-area constraints. Third, we optimize the low-precision dataflow of different LLM modules by applying operator fusion to minimize the overhead of runtime dequantization. Our evaluation on a diverse set of representative LLMs and tasks demonstrates that P3-LLM achieves state-of-the-art quantization accuracy in terms of both KV-cache-only quantization and weight-activation quantization. Combining the proposed quantization scheme with PIM architecture co-design, P3-LLM yields an average of $4.9\times$, $2.0\times$, and $3.4\times$ speedups over the state-of-the-art LLM accelerators HBM-PIM, Ecco, and Pimba, respectively. Our quantization code is available at https://github.com/yc2367/P3-LLM.git
@@ I Introduction
Large language models (LLMs) have revolutionized various machine learning tasks such as text generation [gpt3, llama, mistral, qwen-2p5], image understanding [llama-3.2-vision, qwen-vl], and logical reasoning [deepseek-r1, openai-o3, llama-4]. Performing LLM inference requires the system to effienctly support two distinct stages: prefilling and decoding, each with unique computational characteristics [llm-survey]. During the prefilling stage, the LLM receives the user prompt containing many input tokens and perform general matrix-matrix multiplications (GEMM) to generate the first output token. The GEMM operation can be effectively accelerated by a neural processing unit (NPU) with high computational parallelism [tpuv4i]. Yet during the decoding stage, the LLM produces new tokens in an autoregressive manner, where in every iteration, the LLM only processes one token and performs general matrix-vector multiplications (GEMV) that demand high off-chip memory bandwidth [samsung-aim, sk-aim].
To satisfy the disparate demands of prefilling and decoding stages, heterogeneous LLM acceleration systems based on NPU and DRAM-based processing-in-memory (PIM) have been actively explored by recent literature [neupim, ianus, attacc, sk-pim, samsung-pim, pimba]. An integrated NPU-PIM accelerator leverages the high-performance NPU for compute-intensive GEMM during prefilling, while the high-bandwidth PIM is suitable for memory-intensive GEMV during decoding. Despite PIM’s great promise in accelerating LLM decoding, existing designs for PIM compute units (PCUs) using high-precision (e.g. FP16) arithmetic suffer from significant area and power overhead due to the much lower transistor density in DRAM technology [upmem]. This PCU overhead severely limits the computational throughput of PIM, restricting its practical speedup benefits mainly to single-batch inference on early-generation LLMs, whose multi-head attention and feed-forward layers exhibit an arithmetic intensity of one during decoding [llama, llama-2, gpt3]. Recently, low-batch edge LLM inference has gained much popularity [h2-llm], and state-of-the-art (SoTA) LLMs have adopted grouped-query attention (GQA) with arithmetic intensities larger than one [llama-3p1, llama-3p2, qwen-2p5, mistral]. This trend necessitates new PCU design with higher throughput, while remaining within the area constraints of DRAM technology.
@@
On the algorithm side, SoTA weight-activation quantization methods like QuaRot [quarot] and QServe [qserve] primarily use conventional integer formats, and rely on calibration datasets to approximate the dynamic behavior of activations and KV-cache in a static, offline manner. Our evaluation (see Section VI-B) reveals that these calibration-based integer quantization methods struggle to balance model accuracy due to overfitting. On the hardware side, adopting low-precision PCUs for LLM inference remains an open research problem in the architecture community. While a recent PIM architecture, Pimba [pimba], adopts low-precision arithmetic to improve the area efficiency of PCUs, it only focuses on 8-bit KV-cache quantization with limited memory saving. Furthermore, offloading the self-attention module to low-precision PCUs introduces additional challenges, due to the difficulty of accurately quantizing attention-scores.
To overcome the above limitations, we propose P3-LLM, an NPU-PIM integrated accelerator for low-Precision edge LLM inference. Given the substantial memory and computation demands of LLMs, P3-LLM employs mixed-precision quantization with 4-bit weights and KV-cache to achieve high compression, as well as 8-bit activations and attention scores to reduce computational complexity. Our key algorithmic innovation is an operand-dependent quantization scheme that leverages hybrid numerical formats to aggressively quantize different LLM operands with minimal accuracy loss. For KV-cache, we propose a dynamic, input-aware smoothing strategy that mitigates outliers without the need for a calibration dataset, thus enabling accurate 4-bit integer quantization while eliminating the risk of overfitting. For attention-scores, we introduce an unsigned 8-bit floating-point format with 4-bit exponent and 4-bit mantissa (FP8-S0E4M4) to offer superior numerical fidelity. For activations, we explore the accuracy-efficiency tradeoffs among different 8-bit quantization options, and identity the direct FP8-E4M3 cast as the optimal choice. Our quantization algorithm is further equipped with an NPU-PIM accelerator co-design, which features high-throughput and area-efficient PCUs to flexibly support hybrid numerical formats. Finally, we architect low-precision dataflow of different LLM modules by applying operator fusion to minimize the overhead of runtime dequantization.
@@ II-A Architecture of Large Language Models (LLMs)
The decoder layer serves as the fundamental component in LLMs, consisting of a self-attention module and a multi-layer perceptron (MLP). The self-attention module begins with three linear layers ($W_{Q}$, $W_{K}$, $W_{V}$) to generate query, key, and value vectors, respectively. In recent LLM architectures [llama, mistral, qwen-2p5], $W_{Q}$ and $W_{K}$ are usually followed by rotary position embedding (RoPE) [rope], which encodes positional information into the query and key vectors through matrix rotation. The generated key and value vectors are also cached in off-chip memory for computation reuse during future decoding iterations, and are therefore referred to as the KV-cache. Then, each query and key-value vector is split into $N_{\text{A}}$ and $N_{\text{KV}}$ heads, respectively, where $N_{\text{A}}$ is the number of attention heads and $N_{\text{KV}}$ is the number of key-value heads. For every attention head, the query vectors are multiplied with the transposed key vectors ($Q\cdot K^{\text{T}}$}), followed by a softmax function to calculate the attention-scores ($P$). The attention-scores are then multiplied with the value vectors ($P\cdot V$}), and the results are passed through a linear output-projection layer ($W_{O}$) to produce the attention output states. The MLP module contains three linear layers: gate-projection ($W_{\text{gate}}$), up-projection ($W_{\text{up}}$), and down-projection ($W_{\text{down}}$) to produce the MLP output states.
@@ II-C LLM Quantization
Additionally, recent literature has explored custom numerical formats that can better adapt to the tensor distribution of LLMs [mant, bitmod, blockdialect, nvfp4, fp8]. MANT [mant] and BlockDialect [blockdialect] develop adaptive data types that assign a per-group optimal number format to accommodate diverse data distributions in group-wise quantization. However, selecting the optimal data type for dynamic operands such as activations and KV-cache brings considerable runtime overhead. BitMoD [bitmod] re-purposes the redundant negative zero encoding in the 4-bit floating-point (FP4) format to minimize the weight quantization error, but does not extend its approach to activation and KV-cache quantization. In addition to research efforts, custom data types have been widely adopted in commercial GPUs. For instance, both NVIDIA Hopper and AMD MI300X architectures support the 8-bit floating-point format with two variants: 4-bit exponent 3-bit mantissa (FP8-E4M3) and 5-bit exponent 2-bit mantissa (FP8-E5M2[fp8].
@@ III Motivation
The above observations motivate our exploration of a mixed-precision configuration, namely W4A8KV4P8. In this scheme, weights (W) and KV-cache (KV) are quantized to 4 bits to maximize memory savings, while activations (A) and attention-scores (P) are retained at 8 bits to mitigate accuracy loss. Although existing W4A8KV4 quantization methods keep attention-scores in FP16 due to its small memory impact  [quarot, qserve, mant], we argue that quantizing attention-scores further enhances hardware efficiency by allowing the $P\cdot V$ operation of self-attention to run on low-precision compute units. Unfortunately, the naïve W4A8KV4P8 quantization with the standard integer format results in large perplexity degradation, as highlighted in Fig. 3(b). To address this, P3-LLM proposes an operand-dependent W4A8KV4P8 quantization strategy using hybrid numerical formats (detailed in Section IV). This approach employs different numerical formats for each operand to minimize their quantization error. As a result, P3-LLM achieves better model performance than SoTA W4A8KV4 integer quantization algorithms, QuaRot [quarot] and QServe [qserve].
@@ III-B Limitations of Existing NPU-PIM Systems
Existing PIM solutions mainly focus on accelerating GEMV operations with an arithmetic intensity of one, thus facing challenges in low-batch decoding and GQA. To elucidate such limitations, we conduct roofline analysis on HBM-PIM supporting FP16 arithmetic [samsung-pim], which has $4\times$ higher bandwidth than normal HBM during PIM operations. Fig. 4 demonstrates that the performance benefits of HBM-PIM gradually disappear as the batch size (BS) and GQA group size (G) approach 4, due to limited computational throughput. On the other hand, NPU remains memory-bound even for moderate BS $\geq 16$}, making it highly desirable to increase the throughput of PIM while exploiting its abundant bandwidth. To address this, P3-LLM leverages W4A8KV4P8 quantization to enable cost-effective LLM inference on compact, low-precision hardware. Compared to HBM-PIM, P3-LLM can integrate $4\times$ PCUs under iso-compute-area constraints. Furthermore, as we will discuss in Section V-D, the low-precision PCU enables $2\times$ higher operating frequency than the FP16 PCU, which effectively doubles the peak throughput. Consequently, P3-LLM offers a superior performance roofline with $8\times$ higher throughput over HBM-PIM.
III-C Limitations of Existing Quantization Co-Design
Numerous co-design solutions have been proposed to accelerate quantized LLMs, as summarized in Table I. BitMoD [bitmod] and Oaken [oaken] focus on 4-bit quantization of weights and KV-cache, respectively. But since all other operands remain in FP16, these two approaches lead to moderate memory savings while incurring considerable hardware cost. In contrast, MANT [mant] and Ecco [ecco] achieve greater memory savings via W4A8KV4 quantization. Nevertheless, MANT exhibits noticeable accuracy degradation under this aggressive quantization scheme, while Ecco requires complicated online decompression of quantized operands back to FP16 with low computational efficiency. Pimba [pimba] introduces a low-precision PIM architecture that exploits the 8-bit microscaling format [microscaling] to mitigate the area overhead of PCUs. However, it leaves the linear layer unquantized, and the FP32 accumulation pipeline in the microscaling data path limits the overall hardware efficiency gain. Unlike prior approaches, P3-LLM seeks to strike a balanced trade-off among memory footprint, model accuracy, and hardware efficiency. By assigning a dedicated numerical format and precision to each LLM operand, P3-LLM minimizes both quantization error and memory usage. Regarding hardware support, it proposes an efficient PIM architecture that accommodates hybrid numerical formats with minimal area overhead.
@@ IV-A KV-Cache Quantization
Profiling of KV-cache Distribution. Fig. 5 illustrates the KV-cache distribution from representative layers and heads of Llama-2-7B and Llama-3.1-8B, using Wikitext-2 dataset [wikitext] with a context length of 4K. Fig. 5(a) and (e) show that the value cache has no outlier pattern, whereas Fig. 5(b) and (f) reveal that the key cache contains distinct outlier channels, which pose significant challenges to quantization. These observations for align with previous findings that certain fixed channels in activations exhibit larger outliers [smoothquant, llm-int8]. However, our profiling uncovers an additional important insight: the key cache of Llama-2-7B and Llama-3.1-8B exhibit drastically different patterns after rotary positional embedding (RoPE) is applied. In Llama-2-7B, the post-RoPE key cache has much less structured outlier patterns. In contrast, the key cache of Llama-3.1-8B shows little difference before and after RoPE. The reason is that Llama-3 supports a much longer sequence length than Llama-2 (128K vs. 4K tokens). As described in Section II-A, RoPE encodes positional information into the key cache via matrix rotation  specifically, by multiplying each key vector with a rotation matrix whose angle is proportional to the token’s position index. When the maximum supported sequence length is long (e.g., 128K in Llama-3), the rotation angles are very small for typical input context lengths (e.g., 4K) targeting edge inference scenarios [h2-llm, anda], and RoPE has minimal impact on the key-cache distribution. Given these insights, we adopt pre-RoPE and post-RoPE key-cache quantization for Llama-2 and Llama-3, respectively, to take the advantage of structured outlier channels.
Dynamic Key-Cache Smoothing. To eliminate outlier channels in key cache ($K$), we propose a dynamic input-aware smoothing strategy. Assume $K$ is of size $N_{\text{T}}\times H$, where $N_{T}$ is the input context length and $H$ is the hidden dimension size, the smoothed key cache $K_{S}$ can be expressed as:
@@
Here, every key channel is divided by a smoothing factor, which is the per-channel absolute maximum, to lie in the numerical range of [1,1][-1,1]. Consequently, the outlier channels are effectively suppressed as illustrated in Fig. 5(d) and (h). To apply dynamic smoothing to both prefilling and decoding stages, we calculate the smoothing factors by taking the per-channel absolute maximum of the key cache generated during prefilling. The smoothing factors are then saved and reused in the decoding stage to scale the newly generated key vectors. This approach provides two benefits. First, since every channel stores a smoothing factor shared by all tokens, the additional memory required is inversely proportional to the context length. Given that the context length can easily reach several hundreds, this memory overhead is $1\%$} and negligible. Second, the calculation of smoothing factors only involves the prefilling context and has no impact on the decoding stage. Our profiling on an A6000 GPU running Llama-3.1-8B reveals that this calculation takes <5{<5} millisecond in total across all layers, even under a context length of 32K  far beyond the typical context length in edge LLM inference [h2-llm, anda]. When considering a very stringent service-level objective in chatbot applications, with a 250-millisecond requirement on the time-to-first-token [distserve], the runtime overhead of dynamic smoothing is still $2\%$}.
@@
While per-channel smoothing has been explored by prior quantization works, our proposed strategy has two notable differences. First, prior methods often rely on a calibration dataset to determine the outlier threshold [oaken] or smoothing factor [smoothquant, qserve] offline, which can lead to overfitting. In contrast, our method does not require any calibration dataset. Second, unlike prior works that focus exclusively on either pre-RoPE [kvquant] or post-RoPE [qserve] key-cache quantization. we investigate both approaches by carefully analyzing the impact of RoPE on the resulting key cache distribution of different models. Section V-A discuss the impact of pre-RoPE and post-RoPE quantization on the co-designed dataflow of P3-LLM.
@@
Figure 6: (a) Operator mapping on P3-LLM during decoding. (b) The quantized dataflow of three GEMV operations: weights @@ activations, query @@ key, attention-score @@ value. For clarity, we use ”@@” and $\times$” to denote GEMV and element-wise multiplication, respectively. (c) The architecture of P3-LLM.

### v2

@@ Abstract
The substantial memory bandwidth and computational demands of large language models (LLMs) present critical challenges for efficient inference. To tackle this, the literature has explored heterogeneous systems that combine neural processing units (NPUs) with DRAM-based processing-in-memory (PIM) for LLM acceleration. However, existing high-precision (e.g., FP16) PIM compute units incur significant area and power overhead in DRAM technology, limiting the effective computation throughput. In this paper, we introduce P3-LLM, a novel NPU-PIM integrated accelerator for LLM inference using hybrid numerical formats. Our approach is threefold: First, we propose a flexible mixed-precision quantization scheme, which leverages hybrid numerical formats to quantize different LLM operands with high compression efficiency and minimal accuracy loss. Second, we architect an efficient PIM accelerator for P3-LLM, featuring enhanced compute units to support hybrid numerical formats. Our careful choice of numerical formats allows to co-design low-precision PIM compute units that significantly boost the computation throughput under iso-area constraints. Third, we optimize the low-precision dataflow of different LLM modules by applying operator fusion to minimize the overhead of runtime dequantization. Evaluation on a diverse set of representative LLMs and tasks demonstrates that P3-LLM achieves state-of-the-art accuracy in terms of both KV-cache quantization and weight-activation quantization. Combining the proposed quantization scheme with PIM architecture co-design, P3-LLM yields an average of $4.9\times$, $2.0\times$, and $3.4\times$ speedups over the state-of-the-art LLM accelerators HBM-PIM, Ecco, and Pimba, respectively. Our quantization code is available at https://github.com/yc2367/P3-LLM.git
@@ I Introduction
Large language models (LLMs) have revolutionized various machine learning tasks such as text generation [gpt3, llama, mistral, qwen-2p5], image understanding [llama-3.2-vision, qwen-vl], and logical reasoning [deepseek-r1, openai-o3, llama-4]. Performing LLM inference requires the system to efficiently support two distinct stages: prefilling and decoding, each with unique computational characteristics [llm-survey]. During the prefilling stage, the LLM receives the user prompt containing many input tokens and perform general matrix-matrix multiplications (GEMM) to generate the first output token. The GEMM operation can be effectively accelerated by a neural processing unit (NPU) with high computational parallelism [tpuv4i]. Yet during the decoding stage, the LLM produces new tokens in an autoregressive manner, where in every iteration, the LLM only processes one token and performs general matrix-vector multiplications (GEMV) that demand high off-chip memory bandwidth [samsung-aim, sk-aim].
To satisfy the disparate demands of prefilling and decoding stages, heterogeneous LLM acceleration based on NPU and DRAM-based processing-in-memory (PIM) has been actively explored by recent literature [neupim, ianus, attacc, sk-pim, samsung-pim, pimba]. An integrated NPU-PIM accelerator leverages NPU for compute-intensive GEMM during prefilling, while PIM is suitable for memory-intensive GEMV during decoding. Despite PIM’s great promise in accelerating LLM inference, the PIM compute units (PCUs) suffer from significant area and power overhead due to the much lower transistor density in DRAM technology [upmem]. This PCU overhead severely limits the computation throughput of PIM, restricting its practical speedup mainly to single-batch inference on early-generation LLMs, whose multi-head attention and feed-forward layers exhibit an arithmetic intensity of one during decoding [llama, llama-2, gpt3]. Recently, low-batch edge LLM inference has gained much popularity [h2-llm], and state-of-the-art (SoTA) LLMs have adopted grouped-query attention (GQA) with arithmetic intensities larger than one [llama-3p1, llama-3p2, qwen-2p5, mistral]. This trend necessitates new PCU design with higher throughput, while remaining within the area constraints of DRAM technology.
@@
On the algorithm side, SoTA weight-activation quantization algorithms like QuaRot [quarot] and QoQ [qserve] primarily use the conventional integer format, and rely on calibration datasets to approximate the dynamic behavior of activations in a static offline manner, which struggle to maintain good accuracy for other datasets due to overfitting. On the hardware side, adopting low-precision PCUs for LLM inference remains an open research problem in the architecture community, mainly because of the challenges in balancing memory saving, model accuracy, and hardware efficiency. MANT [mant] and Ecco [ecco] propose aggressive quantization methods and co-designed NPU accelerators, but suffer from low computational efficiency due to the complicated numerical formats and encoding to represent quantized operands. Thus, these proposals cannot be directly applied to PIM with stringent area constraints. While a recent PIM work, Pimba [pimba], adopts the 8-bit microscaling format to improve the area efficiency of PCUs, it overemphasizes preserving model accuracy with 8-bit KV-cache-only quantization, resulting in low memory savings for edge LLM inference. Furthermore, offloading the self-attention module of LLMs to low-precision PCU introduces additional challenges, such as the difficulty of accurately quantizing attention-scores.
To overcome the above limitations, we propose P3-LLM, an NPU-PIM integrated accelerator for low-Precision edge LLM inference. Given the substantial memory and computation demands of LLMs, P3-LLM employs mixed-precision quantization with 4-bit weights and KV-cache to achieve high compression, as well as 8-bit activations and attention scores to reduce computational complexity. Our key algorithmic innovation is an operand-dependent quantization scheme that leverages hybrid numerical formats to aggressively quantize different LLM operands with minimal accuracy loss. For KV-cache, we propose a dynamic, input-aware smoothing strategy that mitigates outliers without calibration datasets, thus enabling accurate and efficient 4-bit integer quantization while preventing overfitting. For attention-scores, we introduce an unsigned 8-bit floating-point format with 4-bit exponent and 4-bit mantissa (FP8-S0E4M4) to offer superior numerical fidelity. For activations, we explore the accuracy-efficiency tradeoffs among different 8-bit quantization options, and identity the direct FP8-E4M3 cast as the optimal choice. Our quantization algorithm is further equipped with a low-precision PIM co-design, which features high-throughput and area-efficient PCUs to flexibly support hybrid numerical formats. Finally, we architect low-precision dataflow of different LLM modules by applying operator fusion to minimize the overhead of runtime dequantization.
@@ II-A Architecture of Large Language Models (LLMs)
The decoder layer serves as the fundamental component in LLMs, consisting of a self-attention module and a multi-layer perceptron (MLP). The self-attention module begins with three linear layers ($W_{Q}$, $W_{K}$, $W_{V}$) to generate query, key, and value vectors, respectively. In recent LLM architectures [llama, mistral, qwen-2p5], $W_{Q}$ and $W_{K}$ are usually followed by rotary position embedding (RoPE) [rope], which encodes positional information into the query and key vectors through matrix rotation. The generated key and value vectors are also cached in off-chip memory for computation reuse during future decoding iterations, and are therefore referred to as the KV-cache. Then, each query and key-value vector is split into $N_{\text{A}}$ and $N_{\text{KV}}$ heads, respectively, where $N_{\text{A}}$ is the number of attention heads and $N_{\text{KV}}$ is the number of key-value heads. For every attention head, the query vectors are multiplied with the transposed key vectors ($Q\cdot K^{\text{T}}$}), followed by a softmax function to calculate the attention-scores ($P$). The attention-scores are then multiplied with the value vectors ($P\cdot V$}), and the results are passed through a linear output-projection layer ($W_{O}$) to produce the attention output states. The MLP module contains three linear projection layers ($W_{\text{gate}},W_{\text{up}},W_{\text{down}}$) to produce the MLP output states.
@@ II-C LLM Quantization
Additionally, recent literature has explored custom numerical formats that can better adapt to the tensor distribution of LLMs [mant, bitmod, blockdialect, fp8, amove]. MANT [mant] and BlockDialect [blockdialect] develop adaptive data types that select an optimal numerical format to quantize each group of operands. However, selecting the optimal format for dynamic operands such as activations and KV-cache brings considerable runtime overhead. BitMoD [bitmod] re-purposes the redundant negative zero encoding in the 4-bit floating-point (FP4) format to minimize the weight quantization error. In addition to research efforts, custom data types have been widely adopted by industry. For instance, the flagship GPUs of NVIDIA and AMD support the 8-bit floating-point format with two variants: 4-bit exponent 3-bit mantissa (FP8-E4M3) and 5-bit exponent 2-bit mantissa (FP8-E5M2[fp8]. The latest NVIDIA Blackwell architecture introduces the NVFP4 format [nvfp4], which enhances the standard FP4 data type with a two-level scaling strategy.
@@ III Motivation
The above observations motivate our mixed-precision quantization scheme, namely W4A8KV4P8. In this scheme, weights (W) and KV-cache (KV) are quantized to 4 bits to maximize memory savings, while activations (A) and attention-scores (P) are retained at 8 bits to mitigate accuracy loss. Although attention-scores account for a small memory footprint, we argue that quantizing attention-scores further enhances hardware efficiency by allowing the $P\cdot V$ operation of the self-attention module to run on low-precision compute units. Unfortunately, the naïve W4A8KV4P8 quantization with the standard integer format results in large perplexity degradation, as highlighted in Fig. 3(b). To address this, P3-LLM proposes an operand-dependent W4A8KV4P8 quantization strategy using hybrid numerical formats (detailed in Section IV). This approach employs different numerical formats for each operand to minimize their quantization error. As a result, P3-LLM achieves better model performance than SoTA W4A8KV4 integer quantization algorithms, QuaRot [quarot] and QServe [qserve].
@@ III-B Limitations of Existing NPU-PIM Systems
Existing PIM solutions mainly focus on accelerating GEMV operations with an arithmetic intensity of one, thus facing challenges in low-batch decoding and GQA. To elucidate such limitations, we conduct roofline analysis on HBM-PIM supporting FP16 arithmetic [samsung-pim], which has $4\times$ higher bandwidth than normal HBM during PIM operations. Fig. 4 demonstrates that the performance benefits of HBM-PIM gradually disappear as the batch size (BS) and GQA group size (G) approach 4, due to limited computation throughput. On the other hand, NPU remains memory-bound even for moderate BS $\geq 16$}, making it highly desirable to increase the throughput of PIM while exploiting its abundant bandwidth. To address this, P3-LLM leverages W4A8KV4P8 quantization to enable cost-effective LLM inference on compact, low-precision hardware. Compared to HBM-PIM, P3-LLM can integrate $4\times$ PCUs under iso-compute-area constraints. Furthermore, as we will discuss in Section V-D, the low-precision PCU enables $2\times$ higher operating frequency than the FP16 PCU, which effectively doubles the peak throughput. Consequently, P3-LLM offers a superior performance roofline with $8\times$ higher throughput over HBM-PIM.
III-C Limitations of Existing LLM Accelerator Co-Design
Numerous co-design solutions have been proposed to accelerate quantized LLMs, as summarized in Table I. BitMoD [bitmod] and Oaken [oaken] focus on 4-bit quantization of weights and KV-cache, respectively. But since all other operands remain in FP16, these two approaches lead to moderate memory savings and hardware efficiency. MANT [mant] and Ecco [ecco] achieve greater memory savings via W4A8KV4 quantization. Nevertheless, MANT exhibits noticeable accuracy drop under this aggressive quantization scheme since it does not address the outlier issue presented in activations and KV-cache. Whereas Ecco requires complicated online decompression of quantized data back to FP16 for computation, leading to considerable area overhead. Thus, although Ecco achieves high accuracy, its idea cannot be applied to designing low-precision PCUs for more efficient LLM decoding. On the other hand, Pimba [pimba] introduces a low-precision PIM architecture that exploits the 8-bit microscaling format [microscaling] to mitigate the area overhead of PCUs. However, it leaves the linear layer unquantized, and the FP32 accumulation pipeline in the microscaling data path limits the overall hardware efficiency gain. Unlike these proposals, P3-LLM seeks to strike a balanced trade-off among memory footprint, model accuracy, and hardware efficiency. By assigning a dedicated precision and efficient numerical format to each LLM operand, P3-LLM minimizes both quantization error and memory usage. Our careful choice of hybrid numerical formats allows to co-design low-precision PIM that significantly boosts the computation throughput under iso-compute-area constraints.
@@ IV-A KV-Cache Quantization
Profiling of KV-cache Distribution. Fig. 5 illustrates the KV-cache distribution from representative layers and heads of Llama-2-7B and Llama-3.1-8B, using Wikitext-2 dataset [wikitext] with a context length of 4K. Fig. 5(a) and (e) show that the value cache has no outlier pattern, whereas Fig. 5(b) and (f) reveal that the key cache contains distinct outlier channels, which pose significant challenges to quantization. These observations for align with previous findings that certain fixed channels in activations exhibit large outliers [smoothquant, llm-int8]. However, our profiling uncovers an additional important insight: the key cache of Llama-2-7B and Llama-3.1-8B exhibit drastically different patterns after rotary positional embedding (RoPE) is applied. In Llama-2-7B, the post-RoPE key cache has much less structured outlier patterns. In contrast, the key cache of Llama-3.1-8B shows little difference before and after RoPE. The reason is that Llama-3 supports a much longer sequence length than Llama-2 (128K vs. 4K tokens). As described in Section II-A, RoPE encodes positional information into the key cache via matrix rotation  specifically, by multiplying each key vector with a rotation matrix whose angle is proportional to the token’s position index. When the maximum supported sequence length is long (e.g., 128K in Llama-3), the rotation angles are very small for typical input context lengths (e.g., 4K) targeting edge inference scenarios [h2-llm, anda], and RoPE has minimal impact on the key-cache distribution. Given these insights, we adopt pre-RoPE and post-RoPE key-cache quantization for Llama-2 and Llama-3, respectively, to take the advantage of structured outlier channels.
Dynamic Key-Cache Smoothing. To eliminate outlier channels in the key cache ($K$), we propose a dynamic input-aware smoothing strategy. Assume $K$ is of size $N_{\text{T}}\times H$, where $N_{T}$ is the input context length and $H$ is the hidden dimension size, the smoothed key cache $K_{S}$ can be expressed as:
@@
Here, every key channel is divided by a smoothing factor, which is the per-channel absolute maximum, to lie in the numerical range of [1,1][-1,1]. Consequently, the outlier channels are effectively suppressed as illustrated in Fig. 5(d, h). To apply dynamic smoothing to both prefilling and decoding stages, we calculate the smoothing factors by taking the per-channel absolute maximum of the key cache generated during prefilling. The smoothing factors are then saved and reused in the decoding stage to scale the newly generated key vectors. This approach provides two benefits. First, since every channel stores a smoothing factor shared by all tokens, the additional memory required is inversely proportional to the context length. Given that the context length can easily reach several hundreds, this memory overhead is $1\%$} and negligible. Second, the calculation of smoothing factors only involves the prefilling context and has no impact on the decoding stage. Our profiling on an A6000 GPU running Llama-3.1-8B reveals that this calculation takes <5{<5} millisecond in total across all layers, even under a context length of 32K  far beyond the typical context length in edge LLM inference [h2-llm, anda]. When considering a very stringent service-level objective in chatbot applications, with a 250-millisecond requirement on the time-to-first-token [distserve], the runtime overhead of dynamic smoothing is still $2\%$}.
@@
While per-channel smoothing has been explored by prior quantization works, our proposed strategy has two notable differences. First, prior methods often rely on a calibration dataset to determine the outlier threshold [oaken] or smoothing factor [smoothquant, qserve] offline, which can lead to overfitting. In contrast, our method does not require any calibration dataset. Second, unlike prior works that focus exclusively on either pre-RoPE [kvquant] or post-RoPE [qserve] key-cache quantization. we investigate both approaches by carefully analyzing the impact of RoPE on the resulting key cache distribution of different models. Section V-B discuss the impact of pre-RoPE and post-RoPE quantization on the co-designed dataflow of P3-LLM.
@@
Figure 6: (a) The architecture of P3-LLM. (b) Operator mapping on P3-LLM during decoding. (c) The quantized dataflow of three GEMV operations: weights @@ activations, query @@ key, attention-score @@ value. For clarity, we use ”@@” and $\times$” to denote GEMV and element-wise multiplication, respectively.